# EDEN: Efficient Dual-Layer Exploration Planning for Fast UAV Autonomous Exploration in Large 3-D Environments

### v1

@@ IV-B Curvature-Penalized Viewpoints Planning
After finding the long-term routing of regions, three viewpoints (an exploring viewpoint, a continuous viewpoint, and a safety viewpoint) in the first region of the routing will be selected for ASEO trajectory planning. The UAV will go to explore the exploring viewpoint (in the explored map) and maintain a speed and an acceleration conducive to explore the continuous viewpoint (in the explored map or unexplored map) if no obstacle is detected, otherwise it can return to the safety viewpoint (in the explored map).
For detailed viewpoint evaluation, existing methods usually use the path distance and turning angle to avoid long traveling distances and sharp turns in a decoupled manner. However, the maximum velocity is limited by the curvature of the trajectory at a point according to the circular motion model. We compute the velocity $v_{s}(\mathbf{p})$ at $\mathbf{p}$ by
@@
where $v_{m}$ is the maximum linear velocity of the UAV, $a_{m}$ is the maximum linear acceleration of the UAV, $K_{c}(\mathbf{p})$ is the curvature of the trajectory at $\mathbf{p}$, and $\lambda_{a}\in(0.0,1.0)$ is an adjustable parameter. Based on this fact, we designed a curvature-penalized heuristic motion cost evaluation method to select the best exploring target and the best continuous target for high-speed exploration.
@@ IV-B1 Curvature-Penalized Exploring Viewpoint and Continuous Viewpoint Planning
First, a Dijkstra search starts from the current UAV position in the explored space is conducted to find the path to each candidate viewpoint in the first region of the routing. Then, for each path to a candidate viewpoint, an anchor point on the path which is $\lambda_{arc}\cdot v_{m}$ ($\lambda_{arc}$ is a tunable parameter) away from the UAV position is selected to formulate a virtual circular arc trajectory. The formulation of the $i$-th virtual circular arc trajectory to the $k$-th candidate viewpoint $\textbf{vp}_{k}=\left\{\mathbf{p}_{k},\psi_{k}\right\}\in\mathbb{R}^{3}\times SO(2)$ is constrained by the current UAV position $\mathbf{p}_{c}\in\mathbb{R}^{3}$, the current UAV speed $\dot{\mathbf{p}}_{c}\in\mathbb{R}^{3}$, and the anchor point $\mathbf{p}_{a,k}\in\mathbb{R}^{3}$. The center point of the circular arc $\mathbf{p}_{ac,k}\in\mathbb{R}^{3}$ and its curvature can be computed by
$\mathbf{p}_{ac,k}=$
[…]
$K_{c,k}=\frac{1}{\|\mathbf{p}_{ac,k}-\mathbf{p}_{c}\|_{2}}.$ (3)
@@
The virtual velocity $v_{as,k}$ to $\textbf{vp}_{k}$ is penalized by the curvature $K_{c,k}$ and is computed by Eq. 2. The motion cost of going to this exploring target is computed by
$\mathrm{CostE}(\mathbf{vp}_{k})=\max\left(\frac{\mathrm{PL}(\mathbf{p}_{k},\mathbf{p}_{c})}{v_{as,k}},\frac{\mathrm{YD}(\psi_{k},\psi_{c})}{\dot{\psi}_{m}}\right),$ (4)
@@
where the function $\mathrm{PL}(\mathbf{p}_{c}$, $\mathbf{p}_{k})$ computes the length of the path from $\mathbf{p}_{c}$ to $\mathbf{p}_{k}$, $\mathrm{YD}(\psi_{k},\psi_{c})$ the yaw difference between the current UAV yaw $\psi_{c}$ and $\psi_{k}$, and $\dot{\psi}_{m}$ is the maximum yaw velocity. Since the terminal velocity at $\mathbf{vp}_{k}$ is unknown, we continue to use the $v_{as,k}$ for computing the continuous motion cost from $\mathbf{vp}_{k}$ to the $j$-th viewpoint $\mathbf{vp}_{j}=\left\{\mathbf{p}_{j},\psi_{j}\right\}\in\mathbb{R}^{3}\times SO(2)$ inside the EROI of $\mathbf{vp}_{k}$
$\mathrm{CostC}(\mathbf{vp}_{j})=\max\left(\frac{\mathrm{PL}(\mathbf{p}_{k},\mathbf{p}_{j})}{v_{as,k}},\frac{\mathrm{YD}(\psi_{j},\psi_{k})}{\dot{\psi}_{m}}\right).$ (5)
Finnally, the gain of this viewpoint-pair is evaluated by
$\mathrm{G}(\mathbf{vp}_{k},\mathbf{vp}_{j})=\mathrm{e}^{-\lambda_{e}\mathrm{CostE}(\mathbf{vp}_{k})}\cdot(1+\mathrm{e}^{-\lambda_{e}\mathrm{CostE}(\mathbf{vp}_{j})}).$ (6)
@@
Here, $\lambda_{e}$ is a tunable parameter. The larger $\lambda_{e}$ is, the more sensitive it is to the motion costs. The viewpoint-pair with the highest gain is selected as the exploring viewpoint $\mathbf{vp}_{e}$ and the continuous viewpoint $\mathbf{vp}_{c}$.
@@
It is worth noting that the voxels on the continuous line between $\mathbf{p}_{i}$ and $\mathbf{p}_{j}$ can not be occupied but free and unknown. This ensures the existence of the possibility for continuous exploration. To efficiently check the feasibilities of the viewpoint pairs, we propose a continuous line-checking method that follows the fast collision checking fashion [36] to avoid recalculation of indexes of voxels at overlapping positions of continuous lines. Specifically, our method generates a table of indexes of voxel that needs to be checked in advance and iterates the table when checking an EROI. Once an iterative voxel is found occupied, its corresponding viewpoint pairs are confirmed infeasible and will not be considered.
@@ IV-B2 Safety Viewpoint Planning
After finding $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$, another viewpoint in the first region of routing will be selected. The safety viewpoint $\mathbf{vp}_{s}$ is the backup target if new obstacles are detected on the trajectory between $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$. To find $\mathbf{vp}_{c}$ efficiently, we directly conduct the Dijkstra algorithm starting from $\mathbf{vp}_{e}$, and the viewpoint that closest to $\mathbf{vp}_{e}$ will be selected as $\mathbf{vp}_{c}$.
@@
Figure 3: An instance of ASEO trajectory. The initial position, velocity, and acceleration of $\mathcal{J}_{e}$ (violet) are fixed with the current UAV position, velocity, and acceleration (yellow). The continuous trajectory $\mathcal{J}_{c}$ (red) ends at $\mathbf{vp}_{c}$. The safety trajectory $\mathcal{J}_{s}$ (red) ends at $\mathbf{vp}_{s}$. The optimizable parameters, including intermediate 4-D points, trajectory durations, velocity, and acceleration at $\mathbf{vp}_{e}$, are illustrated in green.
IV-C ASEO Trajectory Optimization
@@
Inspired by [23], we propose to optimize our 4-D aggressive and safe exploration-oriented (ASEO) trajectory which is based on MINCO trajectory [37]. Different from the 3-D MINCO trajectory, the fourth dimension is the yaw angle, and we optimize it with the position trajectory together.
@@
A 4-D ASEO trajectory $\mathcal{J}(t):\left[t_{s},t_{e}\right]\mapsto\mathbb{R}^{3}\times SO(2)$ ($t_{s}$ and $t_{e}$ are the start time and end time of the trajectory) is paramerized by the initial state $\textbf{S}_{i}=\left\{\mathbf{vp}_{i},\dot{\mathbf{vp}}_{i},\ddot{\mathbf{vp}}_{i}\right\}$, the terminal state $\mathbf{S}_{t}=\left\{\mathbf{vp}_{t},\dot{\mathbf{vp}}_{t},\ddot{\mathbf{vp}}_{t}\right\}$, the $m-1$ intermediate viewpoints $\mathbf{VP}=\left\{\mathbf{vp}_{1},\mathbf{vp}_{2},\cdots,\mathbf{vp}_{m-1}\right\}$, and $m$ durations of trajectory segments $\Delta\mathbf{T}=\left\{\Delta t_{1},\Delta t_{2},\cdots,\Delta t_{m}\right\}$. As shown by Fig. 3, the proposed ASEO trajectory is composed of three connected MINCO trajectories: an exploring trajectory $\mathcal{J}_{e}$ (paramerized by $\textbf{S}_{i,e}$, $\textbf{S}_{t,e}$, $\mathbf{VP}_{e}$, and $\Delta\mathbf{T}_{e}$), a continuous trajectory $\mathcal{J}_{c}$ (paramerized by $\textbf{S}_{i,c}$, $\textbf{S}_{t,c}$, $\mathbf{VP}_{c}$, and $\Delta\mathbf{T}_{c}$), and a safety trajectory $\mathcal{J}_{s}$ (paramerized by $\textbf{S}_{i,s}$, $\textbf{S}_{t,s}$, $\mathbf{VP}_{s}$, and $\Delta\mathbf{T}_{s}$). Some parameters of the ASEO trajectory are fixed. The exploring trajectory starts from the current UAV position $\{\mathbf{p}_{c},\psi_{c}\}$, velocity $\{\dot{\mathbf{p}}_{c},\dot{\psi}_{c}\}$, and acceleration $\{\ddot{\mathbf{p}}_{c},\ddot{\psi}_{c}\}$. The three trajectories are connected at the exploring viewpoint $\mathbf{vp}_{e}$. The continuous trajectory and the safety trajectory end at $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$ correspondingly, their velocity and acceleration are set by zero. We summarize these fixed parameters as:
@@
$\mathbf{vp}_{t,e}=\mathbf{vp}_{i,c}=\mathbf{vp}_{i,s}=\mathbf{vp}_{e},$
$\mathbf{vp}_{t,c}=\mathbf{vp}_{c},\mathbf{vp}_{t,s}=\mathbf{vp}_{s},$
@@
The rest of the parameters are optimizable, and these parameters are shown in Fig. 3 (green). To ensure smoothness, three trajectories share the same velocity and acceleration at the connected point.
The objective function of the ASEO trajectory optimization problem is formulated as follows:
@@
where $j_{e,f},j_{c,f}$, and $j_{s,f}$ are the feasibility costs of $\mathcal{J}_{e},\mathcal{J}_{c}$, and $\mathcal{J}_{s}$ corresponingly. To speed up the continuous exploration, $j_{e,t}$ and $j_{c,t}$ (the time costs of $\mathcal{J}_{e}$ and $\mathcal{J}_{c}$) are optimized. The feasibility cost $j_{\mu,f},\mu\in\{e,c,s\}$, it is computed by
$j_{\mu,f}=\int_{t_{s},\mu}^{t_{e,\mu}}\lambda_{p}\mathrm{Vp}_{\mu}(t))+\lambda_{d}(\mathrm{Vv}_{\mu}(t)+\mathrm{Va}_{\mu}(t))dt,$ (9)
where $\mathrm{Vp}_{\mu}(t)$ calculates the position violation of collision-free corridor constraint, $\mathrm{Vv}_{\mu}(t)$ calculates the linear velocity violation of maximum linear velocity constraint, and $\mathrm{Va}_{\mu}(t)$ calculates the maximum acceleration violation of linear acceleration constraint. The penalty weights are $\lambda_{p}$ and $\lambda_{d}$. For the time cost $j_{\mu,t},\mu\in\{e,c\}$, it is computed by
$j_{\mu,f}=\lambda_{t}\sum\limits_{i=1}^{m_{\mu}}\Delta t_{\mu,i},$ (10)
where $t_{\mu,i}\in\Delta T_{\mu}$, $m_{\mu}$ is the number of trajectory segments, and $\lambda_{t}$ is the penalty weight of time. For details of derivation and optimization, we refer the readers to [23].
After generating the ASEO trajectory, $\mathcal{J}_{e}$ is executed. Once $\mathbf{vp}_{e}$ is explored or obstacles are detected on $\mathcal{J}_{c}$, the total planner will plan again.
V experiments and results
V-A Implementation Details
To evaluate the proposed method, we compare our method against state-of-the-art methods [20, 16] in simulation experiments, and conduct real-world experiments.
For comparative simulations, all the methods are implemented in C++ on a computer that runs the Ubuntu 20.04 LTS operation system, it uses Robot Operation System (ROS) for communication between the planner and simulator. The computer is with an Intel Core i7-12700H, a GeForce 3060 12G, and 16G memory. For a fair comparison, all the methods are simulated on the platform that is provided by [20]. The FOV of the RGB-D camera used in the simulation is $[115,92]^{\circ}$, and its maximum sensing range is $5.0\,m$. As for real-world experiments, the proposed method is deployed on a self-made UAV platform (Fig.6 (a)). The proposed method is run on an on-board computer with an Intel Core i7-1260P. The UAV is equipped with a Livox Mid 360 LiDAR for localization and an Intel Realsense D435i RGB-D camera for exploration. The FOV of the Realsense camera is $[87,58]^{\circ}$, and its maximum sensing range is $5.0\,m$. The parameters of our method are shown in Table. I.

### v2

@@ IV-B Curvature-Penalized Viewpoints Planning
After finding the long-term routing of regions, three viewpoints (an exploring viewpoint, a continuous viewpoint, and a safety viewpoint) in the first region of the routing will be selected for ASEO trajectory planning. The UAV will go to explore the exploring viewpoint (in the explored map) and maintain a speed and an acceleration conducive to exploring the continuous viewpoint (in the explored map or unexplored map) if no obstacle is detected, otherwise it can return to the safety viewpoint (in the explored map).
For detailed viewpoint evaluation, existing methods usually use the path distance and turning angle to avoid long traveling distances and sharp turns in a decoupled manner. However, the maximum velocity is limited by the curvature of the trajectory at a point according to the circular motion model. The maximum velocity $v_{s}(\mathbf{p})$ at $\mathbf{p}$ is computed by
@@
where $v_{m}$ is the maximum linear velocity of the UAV, $a_{m}$ is the maximum linear acceleration of the UAV, $K_{c}(\mathbf{p})$ is the curvature of the trajectory at $\mathbf{p}$, and $\lambda_{a}\in(0.0,1.0)$ is a tunable parameter for reducing sharp turns with high accelerations. In the following viewpoints planning, a smaller $\lambda_{a}$ leads to fewer turns but may miss the opportunities to explore some corner areas. Based on this fact, we designed a curvature-penalized motion score evaluation method to select the best exploring target and the best continuous target for high-speed exploration.
@@ IV-B1 Curvature-Penalized Exploring Viewpoint and Continuous Viewpoint Planning
First, a Dijkstra search, which starts from the current UAV position in the explored space, is conducted to find the path to each candidate viewpoint in the first region of the routing. Then, for each path to a candidate viewpoint, an anchor point on the path, which is $1.0\,s\times v_{m}$ length away from the UAV or the path terminal if the path is shorter than $1.0\,s\times v_{m}$, is selected to formulate a virtual circular arc trajectory. The formulation of a virtual circular arc trajectory to the $j$-th candidate viewpoint $\mathbf{vp}_{j}=\left\{\mathbf{p}_{j},\psi_{j}\right\}\in\mathbb{R}^{3}\times[-\pi,\pi)$ is constrained by the current UAV position $\mathbf{p}_{u}\in\mathbb{R}^{3}$, the current UAV speed $\dot{\mathbf{p}}_{u}\in\mathbb{R}^{3}$, and the anchor point $\mathbf{p}_{a,j}\in\mathbb{R}^{3}$. The curvature $K_{u,j}$ of the virtual circular arc trajectory can be computed by
$\theta=\arccos$ $(\frac{\dot{\mathbf{p}}_{u}^{\mathrm{T}}\cdot(\mathbf{p}_{a,j}-\mathbf{p}_{u})}{\|\dot{\mathbf{p}}_{u}\|_{2}\cdot\|\mathbf{p}_{a,j}-\mathbf{p}_{u}\|_{2}}),$
$K_{u,j}$ $={\color[rgb]{0.0,0.0,0.0}\frac{2\sin(\theta)}{\|\mathbf{p}_{a,j}-\mathbf{p}_{u}\|_{2}}.}$ (3)
@@
The virtual velocity $v_{as,j}$ to $\mathbf{vp}_{j}$ is penalized by the curvature $K_{u,j}$ and is computed by Eq. 2. The motion cost of going to this exploring target is computed by
$\mathrm{CostE}(\mathbf{vp}_{j})=\max\left(\frac{\mathrm{PL}(\mathbf{p}_{j},\mathbf{p}_{u})}{v_{as,j}},\frac{|\mathrm{YD}(\psi_{j},\psi_{u})|}{\dot{\psi}_{m}}\right),$ (4)
@@
where the function $\mathrm{PL}(\mathbf{p}_{j}$, $\mathbf{p}_{u})$ computes the length of the path from $\mathbf{p}_{u}$ to $\mathbf{p}_{j}$, $\mathrm{YD}(\psi_{j},\psi_{u}):\left[-\pi,\pi\right)\times[-\pi,\pi)\mapsto[-\pi,\pi)$ computes the yaw difference between the current UAV yaw $\psi_{u}$ and $\psi_{j}$ by angle unwrapping, and $\dot{\psi}_{m}$ is the maximum yaw velocity. Since the terminal velocity at $\mathbf{vp}_{j}$ is unknown, we continue to use the $v_{as,j}$ for computing the continuous motion cost from $\mathbf{vp}_{j}$ to the $k$-th viewpoint $\mathbf{vp}_{k}=\left\{\mathbf{p}_{k},\psi_{k}\right\}\in\mathbb{R}^{3}\times[-\pi,\pi)$ inside the EROI of $\mathbf{vp}_{j}$
$\mathrm{CostC}(\mathbf{vp}_{k})=\max\left(\frac{\mathrm{PL}(\mathbf{p}_{k},\mathbf{p}_{j})}{v_{as,j}},\frac{|\mathrm{YD}(\psi_{k},\psi_{j})|}{\dot{\psi}_{m}}\right).$ (5)
Finally, the score for evaluating the viewpoint pair is computed by
${\color[rgb]{0.0,0.0,0.0}\mathrm{G}(\mathbf{vp}_{k},\mathbf{vp}_{j})=\mathrm{e}^{-\lambda_{e}\mathrm{CostE}(\mathbf{vp}_{j})}\cdot(1+\mathrm{e}^{-\lambda_{e}\mathrm{CostC}(\mathbf{vp}_{k})}).}$ (6)
@@
Here, $\lambda_{e}$ is a positive tunable parameter. For a larger $\lambda_{e}$, the planner is inclined to choose the exploring viewpoint with a low $\mathrm{CostE}(\mathbf{vp}_{j})$. For a smaller $\lambda_{e}$, the planner is inclined to choose the viewpoint pair with a small total motion cost. The viewpoint-pair with the highest score is selected as the exploring viewpoint $\mathbf{vp}_{e}=\{\mathbf{p}_{e},\psi_{e}\}$ and the continuous viewpoint $\mathbf{vp}_{c}=\{\mathbf{p}_{c},\psi_{c}\}$.
@@
It is worth noting that the voxels on the line between $\mathbf{p}_{j}$ and $\mathbf{p}_{k}$ can not be occupied but are free and unknown. This ensures that $\mathbf{p}_{k}$ can be observed at $\mathbf{p}_{j}$, which makes it possible to explore continuously. To efficiently check the feasibility of the viewpoint pairs, we propose a continuous line-checking method that follows the fast collision checking fashion [41] to avoid recalculation of indices of voxels at overlapping positions of continuous lines. Specifically, our method generates a table of indices of voxels that need to be checked in advance and iterates through the table when checking an EROI. Once an iterative voxel is found occupied, its corresponding viewpoint pairs are confirmed infeasible and will not be considered.
@@ IV-B2 Safety Viewpoint Planning
After finding $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$, another viewpoint in the first region of routing will be selected. The safety viewpoint $\mathbf{vp}_{s}=\{\mathbf{p}_{s},\psi_{s}\}$ will be a feasible backup target, which ensures the UAV being able to return to the safe space, if new obstacles are detected on the trajectory between $\mathbf{vp}_{e}$ and $\mathbf{vp}_{c}$. Without $\mathbf{vp}_{s}$, the UAV could be too fast to brake at $\mathbf{vp}_{e}$. Although we could find the optimal safety viewpoint through solving another ATSP starting from $\mathbf{vp}_{e}$, we directly select the viewpoint closest to $\mathbf{vp}_{e}$ as $\mathbf{vp}_{s}$ to save the computation. This safety viewpoint is found by conducting a Dijkstra search starting from $\mathbf{vp}_{e}$ inside the free space. Once a viewpoint is searched, $\mathbf{vp}_{s}$ is set as this viewpoint, and the Dijkstra search stops.
@@
Figure 3: An instance of ASEO trajectory. The initial position, velocity, and acceleration of $\mathcal{J}_{e}$ (violet) are fixed with the current UAV position, velocity, and acceleration (yellow). The continuous trajectory $\mathcal{J}_{c}$ (red) ends at $\mathbf{vp}_{c}^{\prime}$. The safety trajectory $\mathcal{J}_{s}$ (blue) ends at $\mathbf{vp}_{s}^{\prime}$. The optimizable parameters, including intermediate 4-D points, trajectory durations, velocity, and acceleration at $\mathbf{vp}_{e}^{\prime}$, are illustrated in green.
V ASEO TRAJECTORY OPTIMIZATION
@@
Inspired by the fast and safe trajectory planner (FASTER) [28], we propose to optimize our 4-D aggressive and safe exploration-oriented (ASEO) trajectory, which is based on the MINCO trajectory [42]. Different from the 3-D MINCO trajectory, the fourth dimension is the yaw, and we optimize it with the position trajectory together.
@@
A 4-D MINCO trajectory $\mathcal{J}(t):\left[t_{s},t_{e}\right]\mapsto\mathbb{R}^{4}$ ($t_{s}$ and $t_{e}$ are the start time and end time of the trajectory) is paramerized by the initial state $\textbf{S}_{i}=\left\{\mathbf{vp}_{i},\dot{\mathbf{vp}}_{i},\ddot{\mathbf{vp}}_{i}\right\}$, the terminal state $\mathbf{S}_{t}=\left\{\mathbf{vp}_{t},\dot{\mathbf{vp}}_{t},\ddot{\mathbf{vp}}_{t}\right\}$, the m1m-1 intermediate viewpoints $\mathbf{VP}=\left\{\mathbf{vp}_{1},\mathbf{vp}_{2},\cdots,\mathbf{vp}_{m-1}\right\}$, and $m$ durations of trajectory segments $\Delta\mathbf{T}=\left\{\Delta t_{1},\Delta t_{2},\cdots,\Delta t_{m}\right\}$. As shown by Fig. 3, the proposed ASEO trajectory is composed of three connected 4-D MINCO trajectories: an exploring trajectory $\mathcal{J}_{e}$ (paramerized by $\mathbf{S}_{i,e}$, $\mathbf{S}_{t,e}$, $\mathbf{VP}_{e}$, and $\Delta\mathbf{T}_{e}$), a continuous trajectory $\mathcal{J}_{c}$ (paramerized by $\mathbf{S}_{i,c}$, $\mathbf{S}_{t,c}$, $\mathbf{VP}_{c}$, and $\Delta\mathbf{T}_{c}$), and a safety trajectory $\mathcal{J}_{s}$ (paramerized by $\mathbf{S}_{i,s}$, $\mathbf{S}_{t,s}$, $\mathbf{VP}_{s}$, and $\Delta\mathbf{T}_{s}$). The three trajectories of ASEO are connected at the exploring viewpoint $\mathbf{vp}_{e}$ to ensure the coverage of unknown space, which is different from FASTER [28], whose connecting point of its “Whole” trajectory and “Safety” trajectory can be any feasible point on the “Whole” trajectory. Some parameters of the ASEO trajectory are fixed. The initial state of the exploring trajectory $\mathbf{S}_{i,e}=\left\{\mathbf{vp}_{i,e},\dot{\mathbf{vp}}_{i,e},\ddot{\mathbf{vp}}_{i,e}\right\}$ is aligned with the UAV position $\{\mathbf{p}_{c},\psi_{c}\}$, velocity $\{\dot{\mathbf{p}}_{c},\dot{\psi}_{c}\}$, and acceleration $\{\ddot{\mathbf{p}}_{c},\ddot{\psi}_{c}\}$. $\mathcal{J}_{e}$’s terminal pose $\mathbf{vp}_{t,e}$, $\mathcal{J}_{c}$’s initial pose $\mathbf{vp}_{i,c}$, and $\mathcal{J}_{s}$’s initial pose $\mathbf{vp}_{i,s}$ are fixed to the exploring viewpoint. $\mathcal{J}_{c}$’s terminal pose $\mathbf{vp}_{t,c}$ and $\mathcal{J}_{s}$’s terminal pose $\mathbf{vp}_{t,s}$ are fixed to the continuous viewpoint and the safety viewpoint correspondingly, their terminal velocity and acceleration are set to zero. We summarize these fixed parameters as:
@@
$\mathbf{vp}_{t,e}=\mathbf{vp}_{i,c}=\mathbf{vp}_{i,s}=\mathbf{vp}_{e}^{\prime}=\{\mathbf{p}_{e},\psi_{e}^{\prime}\},$
$\mathbf{vp}_{t,c}=\mathbf{vp}_{c}^{\prime}=\{\mathbf{p}_{c},\psi_{c}^{\prime}\},\mathbf{vp}_{t,s}=\mathbf{vp}_{s}^{\prime}=\{\mathbf{p}_{s},\psi_{s}^{\prime}\},$
@@
The rest of the parameters are optimizable, and these parameters are shown in Fig. 3 (green). To ensure smoothness, three trajectories share the same velocity and acceleration at the connected point. It is worth noting that we use $\psi_{e}^{\prime}$, $\psi_{c}^{\prime}$, $\psi_{s}^{\prime}\in\mathbb{R}$ instead of $\psi_{e}$, $\psi_{c}$, $\psi_{s}\in[-\pi,\pi)$ to avoid the rotations greater than $\pi$. Specifically, $\psi_{e}^{\prime}$, $\psi_{c}^{\prime}$, and $\psi_{s}^{\prime}$ are calculated by $\psi_{e}^{\prime}=\mathrm{YD}(\psi_{e},\psi_{u})+\psi_{u}$, $\psi_{c}^{\prime}=\mathrm{YD}(\psi_{c},\psi_{e})+\psi_{e}^{\prime}$, and $\psi_{s}^{\prime}=\mathrm{YD}(\psi_{s},\psi_{e})+\psi_{e}^{\prime}$. The intermediate yaws of $\mathcal{J}_{e}$ are bounded within $\left[\min(\psi_{e}^{\prime},\psi_{u})-\pi,\max(\psi_{e}^{\prime},\psi_{u})+\pi\right]$, the intermediate yaws of $\mathcal{J}_{c}$ are bounded within $\left[\min(\psi_{c}^{\prime},\psi_{e}^{\prime})-\pi,\max(\psi_{c}^{\prime},\psi_{e}^{\prime})+\pi\right]$, and the intermediate yaws of $\mathcal{J}_{s}$ are bounded within $\left[\min(\psi_{s}^{\prime},\psi_{e}^{\prime})-\pi,\max(\psi_{s}^{\prime},\psi_{e}^{\prime})+\pi\right]$.
Unlike FASTER, which optimizes different trajectories in a decoupled manner, we optimize the three trajectories of the ASEO trajectory simultaneously to ensure that each trajectory is feasible. The objective function of the ASEO trajectory optimization problem is formulated as follows:
@@
where $j_{e,f},j_{c,f}$, and $j_{s,f}$ are the feasibility costs of $\mathcal{J}_{e},\mathcal{J}_{c}$, and $\mathcal{J}_{s}$ correspondingly. To speed up the continuous exploration, $j_{e,t}$ and $j_{c,t}$ (the time costs of $\mathcal{J}_{e}$ and $\mathcal{J}_{c}$) are optimized. The feasibility cost $j_{\mu,f},\mu\in\{e,c,s\}$ is computed by
$j_{\mu,f}$ $=\int_{t_{s},\mu}^{t_{e,\mu}}\lambda_{p}\mathrm{Vp}_{\mu}(t)+\lambda_{d}(\mathrm{Vv}_{\mu}(t)+\mathrm{Va}_{\mu}(t)$
$+\mathrm{V\dot{\psi}}_{\mu}(t)+\mathrm{V\ddot{\psi}}_{\mu}(t))dt,$ (9)
where $\mathrm{Vp}_{\mu}(t)$ calculates the position violation of collision-free corridor constraint, and $\mathrm{Vv}_{\mu}(t)$, $\mathrm{Va}_{\mu}(t)$, $\mathrm{V\dot{\psi}}_{\mu}(t)$, and $\mathrm{V\ddot{\psi}}_{\mu}(t)$ calculate the violations of linear velocity, linear acceleration, yaw velocity, and yaw acceleration, respectively. The penalty weights are $\lambda_{p}$ and $\lambda_{d}$. For the time cost $j_{\mu,t},\mu\in\{e,c\}$, it is computed by
${\color[rgb]{0.0,0.0,0.0}j_{\mu,t}=\lambda_{t}\sum\limits_{i=1}^{m_{\mu}}\Delta t_{\mu,i},}$ (10)
where $\Delta t_{\mu,i}\in\Delta\mathbf{T}_{\mu}$, $m_{\mu}$ is the number of trajectory segments, and $\lambda_{t}$ is the penalty weight of time. For details of derivation and optimization, we refer the readers to [42].
After generating the ASEO trajectory, $\mathcal{J}_{e}$ is executed. Our exploration planner replans (replans viewpoints and regenerates ASEO trajectory) when it encounters the following conditions: (1) the distance from an obstacle to a point on $\mathcal{J}_{e}$, $\mathcal{J}_{c}$, or $\mathcal{J}_{s}$ is shorter than the radius of the UAV $r_{u}$; (2) $\mathbf{vp}_{e}$ does not cover any unknown space or the UAV is about to reach $\mathbf{vp}_{e}$ ($\|\mathbf{p}_{u}-\mathbf{p}_{e}\|_{2}<d_{th}$ and $|\mathrm{YD}(\psi_{e},\psi_{u})|<\psi_{th}$, where $d_{th}$ is the distance threshold and $\psi_{th}$ is the yaw threshold); (3) the UAV has executed $\mathcal{J}_{e}$ for a replanning period $t_{r}$ or $\mathcal{J}_{e}$ will expire within a replanning offset time $t_{o}$.
VI EXPERIMENTS and RESULTS
VI-A Implementation Details
To evaluate the proposed method, we compare our method against state-of-the-art methods [22, 19] in simulation experiments and conduct real-world experiments.
For comparative simulations, all the methods are implemented in C++ on a computer that runs the Ubuntu 20.04 LTS operation system, which uses Robot Operation System (ROS) for communication between the planner and simulator. The computer has an Intel Core i7-12700H, a GeForce 3060 12G, and 16G memory. For a fair comparison, all the methods are simulated on the platform that is provided by [22]. The FOV of the RGB-D camera used in the simulation is $115,92]^{\circ}$, and its maximum sensing range is 5.0m5.0\,m. As for real-world experiments, the proposed method is deployed on a self-made UAV platform (Fig.6 (a)). The proposed method is run on an on-board computer with an Intel Core i7-1260P. The UAV is equipped with a Livox Mid 360 LiDAR for localization (using Point-LIO [43]) and an Intel Realsense D435i RGB-D camera for exploration. The FOV of the Realsense camera is $87,58]^{\circ}$, and its maximum sensing range is 5.0m5.0\,m. The parameters of our method are shown in Table. I.